### v1

# Alphazero-like Tree-Search can guide large language model decoding and training

Abstract
Large language models (LLMs) typically employ sampling or beam search, accompanied by prompts such as Chain-of-Thought (CoT), to boost reasoning and decoding ability. Recent work like Tree-of-Thought (ToT) and Reasoning via Planning (RAP) aim to augment the reasoning capabilities of LLMs by utilizing tree-search algorithms to guide multi-step reasoning. These methods mainly focus on LLMs’ reasoning ability during inference and heavily rely on human-designed prompts to activate LLM as a value function, which lacks general applicability and scalability. To address these limitations, we present an AlphaZero-like tree-search learning framework for LLMs (termed TS-LLM), systematically illustrating how tree-search with a learned value function can guide LLMs’ decoding ability. TS-LLM distinguishes itself in two key ways: (1) Leveraging a learned value function, our approach can be generally applied to different tasks beyond reasoning (such as RLHF alignment), and LLMs of any size, without prompting advanced, large-scale models. (2) It can guide LLM’s decoding during both inference and training. Empirical evaluations across reasoning, planning, and RLHF alignment tasks validate the effectiveness of TS-LLM, even on trees with a depth of 64.
1 Introduction
Pretrained autoregressively on extensive corpora, large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023a) have demonstrated their potential in a wide range of natural language tasks. A plethora of recent studies have concentrated on improving LLMs task-solving capability, including curation of larger and higher-quality general or domain-specific data (Touvron et al., 2023a; Zhou et al., 2023; Gunasekar et al., 2023; Feng et al., 2023; Taylor et al., 2022), more sophisticated prompt design (Wei et al., 2022; Zhou et al., 2022; Creswell et al., 2022), or better training algorithms with Supervised Learning or Reinforcement Learning (RL) (Dong et al., 2023; Gulcehre et al., 2023; Rafailov et al., 2023). When training LLMs with RL, LLMs’ generation can be naturally formulated as a Markov Decision Process (MDP) and optimized with specific objectives. Following this formulation, ChatGPT (Ouyang et al., 2022) emerges as a notable success, optimizing LLMs to align human preference by leveraging RL from Human Feedback (RLHF) (Christiano et al., 2017).
Based on this formulation, LLMs’ generation can be further guided with planning algorithms such as tree search. The success of AlphaZero (Silver et al., 2017b) has strongly proved the effectiveness of tree search algorithms, especially Monte Carlo Tree Search (MCTS) (Kocsis & Szepesvári, 2006; Coulom, 2006), in solving large-scale MDPs, including Atari and Go. It attracts recent efforts on the potential of leveraging these principled planning approaches to improve LLMs’ reasoning ability.
Preliminary work in this field includes Tree-of-Thought (ToT) (Yao et al., 2023; Long, 2023) with depth/breadth-first search and Reasoning-via-Planing (RAP) (Hao et al., 2023) with MCTS. They successfully demonstrated a performance boost of searching on trees expanded by LLM through self-evaluation. Despite the strides, current methods come with distinct limitations. Firstly, these approaches mainly focus on enhancing LLMs’ reasoning ability, lacking general applicability in different kinds of tasks, such as RLHF alignment. Secondly, as a model-based algorithm, the crucial reward or value function of tree-search algorithms is mainly obtained by prompting LLMs for step-level or trajectory-level reasoning. As a result, such algorithms lack their applicability and heavily rely on both well-designed prompts and large-scale LMs, such as GPT4 or LLaMA-33B. Only LLMs in such scale can provide reliably innate critic ability for reward and value estimation.
To address these problems, we introduce tree-search enhanced LLM (TS-LLM), an AlphaZero-like framework that utilizes tree-search to improve LLMs’ performance on general natural language tasks. TS-LLM extends previous work to AlphaZero-like deep tree-search with a learned value function and can guide LLM both during inference and training. Compared with previous work, TS-LLM has the following two new features:
  • TS-LLM offers a generally applicable and scalable pipeline. It is generally applicable: (1) TS-LLM can be versatile to different tasks beyond reasoning. It can even solve RLHF alignment task in our experiment. (2) With a learned value function, TS-LLM can be applied to LLMs of any size. It does not require any well-designed prompts or advanced, large-scale LLMs. Our experiments show that TS-LLM can work for LLMs ranging from as small as 125M to as large as 7B. TS-LLM is also scalable: TS-LLM can conduct deep tree search, extending tree-search for LLM generation up to a depth of 64. This is far beyond 10 in ToT and 7 in RAP.
  • TS-LLM can potentially serve as a new LLM training paradigm beyond inference decoding. By treating the tree-search operation as a policy improvement operator, we can conduct iterative processes of policy improvement and evaluation, to further train the LLM.
Furthermore, we present an in-depth analysis of the core design elements in TS-LLM, delving into the features, advantages, and limitations of different variations. We also offer a novel and reasonable evaluation metric to fairly compare tree-search algorithms with other baselines. Through comprehensive empirical evaluations on reasoning, planning, and RLHF alignment tasks, we benchmark the performance of various TS-LLM  algorithm variants and underscore their superior capabilities. This showcases TS-LLM’s potential as a universal framework to guide LLM decoding and training.
2 Related work
Multistep Reasoning in LLMs Multistep reasoning in language models has been widely studied, from improving the base model (Chung et al., 2022; Fu et al., 2023; Lewkowycz et al., 2022) to prompting LLMs step by step (Kojima et al., 2022; Wei et al., 2022; Wang et al., 2022; Zhou et al., 2022). Besides, a more relevant series of work has focused on enhancing reasoning through evaluations, including learned reward models (Uesato et al., 2022; Lightman et al., 2023) and self-evaluation (Shinn et al., 2023; Madaan et al., 2023). In this work, we apply evaluations to multistep reasoning tasks and token-level RLHF alignment tasks by learning a value function and reward model under the setting of a multistep decision-making process.
Search guided reasoning in LLMs While most CoT approaches have used a linear reasoning structure, recent efforts have been made to investigate non-linear reasoning structures such as trees (Jung et al., 2022; Zhu et al., 2022). More recently, various approaches for searching on trees have been applied to find better reasoning paths, e.g. beam search in Xie et al. (2023), depth-/breadth-first search in Yao et al. (2023) and Monte-Carlo Tree Search in (Hao et al., 2023). Compared with these methods, TS-LLM is a tree search guided LLM decoding and training framework with a learned value function, which is more generally applicable to both reasoning tasks and other scenarios like RLHF alignment tasks. Due to the limit of space, we leave a comprehensive comparison in Appendix A.
Finetuning LLMs with Augmentation Recent efforts have also been made to improve LLMs with augmented data. Rejection sampling is a simple and effective approach for finetuning data augmentation to improve LLMs’ ability on single/multiple task(s) such as multistep reasoning (Yuan et al., 2023a; Zelikman et al., 2022) and alignment with human preference (Dong et al., 2023; Bai et al., 2022). Given an augmented dataset, reinforcement learning approaches have been also used to finetune LLMs (Gulcehre et al., 2023; Luo et al., 2023). Compared to previous works, TS-LLM leverages tree search as a policy improvement operator to generate augmented samples to train both the LLMs and the value function.
Figure 1: Left: Sentence/Token-level node on Game24. Right: Overall framework: TS-LLM is a generally applicable and scalable pipeline with multiple choices of tree search approaches and search aggregation methods. Firstly, with LLM finetuned on the task training set, TS-LLM generates trajectories to learn value/ORM. By treating tree search as a policy improvement operator, it iteratively augments the training set and further improves the LLM and value/ORM.
3 Enhancing Large language model with Tree Search
In this section, we propose a versatile tree-search framework TS-LLM for guiding LLMs decoding and training. TS-LLM is summarized in the right part of Fig.1, and we conduct a systematic and comprehensive discussion about its key components. We start with the problem formulation with different tree/node designs, illustrating how LLMs’ generation can be framed into a tree-search pipeline. Secondly, we show how tree-search can help with LLM inference, including choices of specific tree-search algorithms, value-function training techniques, and search aggregation methods. Building upon tree-search enhanced inference, we present how tree-search can serve as a general policy improvement operator for guiding LLMs training. Lastly, we discuss the tree-search algorithm’s computation burdens for a fair comparison.
3.1 problem formulation
We formulate the language generation process as a multi-step Markov Decision Process (MDP). The particularity of natural language tasks is that both the action and state are in language space. LLMs can serve as a policy $\pi_{\theta}$ that samples sequences of tokens as actions. Assuming the length of the output sequence and input prompt are $T$ and $L$ respectively, the probability for an LLM policy to produce an output sequence $\mathbf{y}=(y_{0},y_{1},\ldots,y_{T-1})$ conditioned on a prompt (input prefix) $\mathbf{x}=(x_{0},x_{1},\ldots,x_{L-1})$ is: $\pi_{\theta}(\mathbf{y}|\mathbf{x})=\prod_{t=0}^{T-1}\pi_{\theta}(y_{t}|\mathbf{x}_{0:L-1},\mathbf{y}_{0:t-1}).$
For a given natural language task, we can define a reward function $R(y_{t}|\mathbf{x}_{0:L-1},\mathbf{y}_{0:t-1})$ as the task performance feedback for intermediate generation $y_{t}$ at timestep $t$. Due to the lack of large-scale and high-quality intermediate reward labels for general tasks, it is usually a sparse reward setting where any intermediate reward from the first $T-1$ timestep is zero except the last $T$-th step. A typical case can be RLHF alignment task, where LLM can only receive the reward signal after it completes the full generation. Following the same logic, $\mathbf{y}$ can also be viewed as a sequence of sentences.
Given the problem formulation above, we successfully transfer the problem of better generation to optimization for higher cumulative reward. In this paper, we focus on how we can optimize it with tree-search algorithms. A specific natural language task typically predefines the state space (with language) and reward function (with task objective/metrics). What remains is the definition of action space, or in the context of tree-search algorithm, the action node design problem.
Tree search algorithms have validated their effectiveness for different action spaces in traditional RL research, including discrete action space (Silver et al., 2017a; Schrittwieser et al., 2020) and continuous action space (Hubert et al., 2021). For conducting tree-search on LLMs, we categorize the action space design into the following two aspects:
  • Sentence-level action nodes: For the tasks that have a step/sentence-level structure(e.g. chain-of-thought reasoning), it is natural to treat each thought as a sentence-level action node. Fig. 1 shows an example of sentence-level action nodes. This is also the technique adopted by ToT (Yao et al., 2023) and RAP (Hao et al., 2023). For each non-terminal node, the search tree is expanded by sampling $m$ possible subsequent intermediate steps and dropping the duplicated generations.
  • Token-level action nodes: Analogous to tree-search in discrete action space MDP, we can treat each token as a discrete action for LLM policy and the tree search can be conducted in token-level. We demonstrate an example of a Token-level action node in Fig. 1. For those tasks in which intermediate steps are not explicitly defined(e.g. RLHF alignment), splitting an output sequence into tokens might be a good choice.
Both action space designs have their own advantages and limitations. By splitting the generation into sentences, sentence-level action nodes provide a relatively shallow tree, simplifying the tree-search process. However, the large sample space of sentence-level generation makes full enumeration of all possible sentences infeasible. We have to subsample $k$ nodes during the expansion, similar to the idea of Sampled MuZero(Hubert et al., 2021). Such subsampling results in the gap, determined by $k$, between the tree-search space and the LLM generation space. In addition, sentence-level action node often brings extra computation burdens, which will be further discussed in Section 3.4. For token-level action nodes, though it can get rid of the search space discrepancy and extra computational burdens, it greatly increases the depth of the tree, making tree-search more challenging.
3.2 Guiding LLM inference decoding with Tree Search
One of the benefits of tree-search algorithms is that they can optimize the cumulative reward by mere search, without any gradient calculation or update. Tree-search algorithms can enhance LLM generation without any further finetuning. In this section, given a fixed LLM policy, we present the full pipeline to illustrate how to guide LLM inference with tree search approaches.
3.2.1 Learning a Value Function
As a model-based RL algorithm, tree search typically requires three models, including dynamic transition model $g$, reward model $\hat{r}$(Schrittwieser et al., 2020), and value function $v$ Silver et al. (2017a). The dynamic transition in the LLM generation, given state $s$ and action $a$, is known as: $g\big{(}s=(\mathbf{x}_{0:L-1},\mathbf{y}_{0:t-1}),a=y_{t}\big{)}=(\mathbf{x}_{0:L-1},\mathbf{y}_{0:t})$. The value function $v$ and reward model $\hat{r}$ are the main issues. ToT and RAP obtain these two models by prompting advanced LLMs, such as GPT-4 or LLaMA-33B, because only LLMs of such scale can serve as sufficiently reliable models.
To make the tree search algorithm generally applicable, our method leverages a learned value function $v_{\theta}(s)$ conditioned on state $s$ and a learned final-step outcome reward model (ORM) $\hat{r}_{\theta}$ (Uesato et al., 2022) since most tasks can be formulated as sparse-reward problems. We utilize a shared value network and reward model, whose structure is a decoder-only transformer with an MLP to output a scalar on each position of the input tokens. For a sentence-level expanded intermediate step $s_{t}$, we use the prediction scalar at the last token as its value prediction $v_{\theta}(s_{t})$. The final reward can be obtained at the last token when feeding the full sentences $(\mathbf{x}_{0:L-1},\mathbf{y}_{0:T-1})$ into the model.
Therefore, we use language model $\pi_{\theta}$ as the policy to sample generations using the task training dataset. With true label or a given reward function in training data, a set of sampled tuple $\mathcal{D}_{train}=\{(\mathbf{x}^{k},\mathbf{y}^{k},r^{k})\}_{k=1}^{N}$ of size $N$ can be obtained, where $\mathbf{x}^{k}$ is the input text, $\mathbf{y}^{k}=s^{k}_{0:T^{k}-1}$ is the output text of $T^{k}$ steps and $r^{k}=R(\mathbf{y}^{k}|\mathbf{x}^{k})$ is the ground-truth reward. Similar to the critic training in most RL algorithms, we construct the value target $z^{k}_{t}$ by TD-$\lambda$ or MC estimate on each single step $t$. The value network is optimized by mean squared error:
$L(\theta)=\displaystyle\mathbb{E}_{\mathcal{D}}\left[\sum\limits_{t=0}^{T^{k}-1}||v_{\theta}(s^{k}_{0:t}|\mathbf{x}^{k})-z^{k}_{t}||_{2}^{2}\right].$ (1)
The ORM $\hat{r}_{\theta}(\mathbf{y}_{0:T-1}|\mathbf{x}_{0:L-1})$ is learned with the same objective. Training an accurate value function and ORM is quite crucial for the tree-search process as they provide the main guidance during the search. We will further illustrate how different training conditions influence the learned value function, ORM, and corresponding tree-search performance in our experiment section.
3.2.2 Tree Search Algorithms
Given a learned value function, in this section, we present five types of tree-search algorithms. Due to the space limitation, we leave the background and preliminaries in Appendix C.
MCTS with value function approximation (named as MCTS-$\alpha$): This is the MCTS variants utilized in AlphaZero (Silver et al., 2017a). Starting from the initial state, we choose the node of state $s_{t}$ as the root node and do $n$ times of search simulations consisting of select, expand and evaluate and backup, where the leaf node value evaluated by the learned value function will be backpropagated to all its ancestor nodes. After the search, we choose an action proportional to the root node’s exponentiated visit count, i.e. $a\sim\frac{N(s_{t},a)^{1/\tau}}{\sum_{b}N(s_{t},b)^{1/\tau}}$, and move to the corresponding next state. The above iteration will be repeated until finished. MCTS-$\alpha$ has two main features. Firstly, MCTS-$\alpha$ cannot trace back to its previous states once it takes an action. So it cannot restart the search from the initial state unless multiple searches are conducted which will be discussed in Section 3.2.3. Secondly, MCTS-$\alpha$ utilizes a value function so it can conduct the backward operation during the intermediate steps, without the need to complete the full generation to obtain a Monte-Carlo estimate.
MCTS: This approach was adopted in RAP (Hao et al., 2023), which refers to classic MCTS(Kocsis & Szepesvári, 2006). Specifically, the process starts from the initial state node and then select until finding a leaf node, if this node is non-terminal, expand it and repeat the previous select and expand steps. If the process encounters a terminal node, its value will be evaluated and backup to the whole path. In contrast to MCTS-$\alpha$, it only back-propagates the value on the terminal nodes, relying on a Monte-Carlo estimate of value, and it always starts searching from the initial state node.
MCTS-Rollout: Combining the features from the two algorithms above, we propose a new variant MCTS-Rollout for tree search. Similar to MCTS, MCTS-Rollout always starts from the initial state node. It further does the search simulations analogous to MCTS-$\alpha$, and the backup process can happen in the intermediate step with value function. It repeats the operations above until the process finds $k$ complete answers or reaches the computation limit (e.g. maximum number of tokens.)
Breadth-first/Depth-first Search (BFS/DFS): These two search algorithms were adopted in ToT (Yao et al., 2023). The core idea is to utilize the value function to prune the tree for efficient search, while such pruning happens in tree breadth and tree depth respectively. BFS can be regarded as a beam-search with cumulative reward as the objective. In BFS, nodes with $k$-largest values are maintained and expanded in each step. DFS chooses an unvisited node with the highest value until it finds a complete path. To improve the search efficiency, DFS prunes trees by either dropping the nodes whose value is below a predefined threshold or dropping the nodes of the $n$-lowest value.
3.2.3 Multiple Search and Search Aggregation
Inspired by Wang et al. (2022) and Uesato et al. (2022) that LLM can improve its performance on reasoning tasks by sampling multiple times and aggregating the candidates, TS-LLM also have the potential to aggregate answers generated by multiple tree searches. There are mainly two choices:
Intra-tree Search. Intra-tree search conducts multiple tree searches on exactly the same tree. Such a method is computationally efficient as the search tree can be reused multiple times. However, the diversity of multiple generations might decrease because the former tree search might influence the latter tree searches. Also, the search space is limited by the sentence-level actions which are previously expanded in the tree and are fixed across multiple tree searches.
Inter-tree Search. Inter-tree Search builds a new tree for each new search, which increases the diversity of the search space, with extra computation burdens proportional to the search times.
Our next step is to aggregate these search results to obtain the final answer. With a learned ORM, we mainly consider the following three different aggregation methods:
Majority Vote. Wang et al. (2022) aggregates the answer $f$ by using majority vote over multiple random generation $y_{i}$: $f^{*}=\arg\max_{f}\sum\nolimits_{\mathbf{y}^{k}}\mathbf{1}_{\text{final\_ans}(\mathbf{y}^{k})=f}$, where $\mathbf{1}$ is the indicator function.
ORM-Max. Given an outcome reward model, the aggregation can choose the answer $f$ with maximum final reward, $f^{*}=\text{final\_ans}(\arg\max_{\mathbf{y}^{k}}\hat{r}_{\theta}(\mathbf{y}^{k}|\mathbf{x}^{k}))$.
ORM-Vote. Given an outcome reward model, the aggregation can choose the answer $f$ with the sum of reward, namely $f^{*}=\arg\max_{f}\sum\nolimits_{\mathbf{y}^{k};\text{final\_ans}(\mathbf{y}^{k})=f}\hat{r}_{\theta}(\mathbf{y}^{k}|\mathbf{x}^{k})$
3.3 Enhancing LLM Training with Tree Search
By plugging in the tree search operation, LLM can decode better compared to its original decoding performance during inference, which in other words, is a policy improvement process. Based on this, a new training/finetuning paradigm can be proposed, consisting of the following iterative process:
Policy improvement: We firstly generate tree-search trajectories based on $\pi_{\theta_{\text{old}}}$, $v_{\theta_{\text{old}}}$, and $\hat{r}_{\theta_{\text{old}}}$. By imitating new trajectories, LLM policy can be further improved to $\pi_{\theta_{\text{new}}}$.
Policy evaluation: Train value function $v_{\theta_{\text{new}}}$ and ORM $\hat{r}_{\theta_{\text{new}}}$ over the new sampled trajectories.
Such iterative process belongs to generalized policy iteration (Sutton & Barto, 2018), which is also the procedure used in AlphaZero’s training. In our case, the training process involves finetuning three networks on the tree-search augmented training dataset.: (1) Policy network $\pi_{\theta}$: Use cross-entropy loss with trajectories’ tokens as target (2) Value network $v_{\theta}$: Mean squared error loss with trajectories’ temporal difference (TD) or Monte-Carlo (MC) based value estimation as target, and (3) ORM $\hat{r_{\theta}}$: Mean squared error loss with trajectories’ final reward as target.
3.4 Tree Search’s extra Computation Burdens
Tree-search algorithms will inevitably bring in additional computation burdens, especially in the node expansion phase for calculating legal child nodes and their corresponding value. Prior methodologies, such as ToT and RAP, tend to benchmark their performance against baseline algorithms using an equivalent number of generation paths (named Path@N). This approach overlooks the additional computational demands of the tree-search process and makes the comparison unfair.
In the setting of sentence-level action nodes, such computational demands are substantial. This is primarily because expanding a tree requires the generation of multiple sentences from a single parent node. In contrast, a token-level action node can alleviate the computation demand for all child nodes. A single forward inference can directly yield a probability distribution for all tokens so its search space and computational overhead align closely with the random sampling strategy typically used in LLM generation. Despite this, additional exploration steps (especially in Alphazero-like MCTS search) still introduce further computation. To enable a more fair comparison between tree-search and other baselines, we also should monitor the number of tokens generated for node expansion in our experiments. This helps us juxtapose the performance of different algorithms when operating under comparable token generation conditions.
4 Experiments
In this section, we first introduce detailed experimental setups in Section 4.1. In Section 4.2, we comprehensively present features of TS-LLM by answering six questions. Refer to Appendix D and E for more settings, implementations, results, and visualizations. Our code is open-sourced at https://github.com/waterhorse1/LLM_Tree_Search.
4.1 Experiment Setups
Table 1: Task setups
Task Category Tree Max width Tree Max depth Node Train/test size
GSM8k Mathematical Reasoning 6 8 Sentence 7.5k / 1.3k
Game24 Mathematical Planning 20 4 Sentence 1.0k / 0.3k
PrOntoQA Logical Reasoning 6 15 Sentence 4.5k / 0.5k
RLHF Alignment 50 64 Token 30k / 3k
Task setups. For a given MDP, the nature of the search space is primarily characterized by two dimensions: depth and width. To showcase the efficacy of tree-search algorithms across varied search spaces, we evaluate all algorithms on four tasks with different search widths and depths, including the mathematical reasoning task GSM8k (Cobbe et al., 2021), mathematical planning task Game24 (Yao et al., 2023), logical reasoning task PrOntoQA (Saparov & He, 2022), and RLHF alignment task using synthetic RLHF data (Dahoas, ). The specific task statistics and search space hyperparameters are listed in Table 1. These hyperparameters, especially max search width and search depth, are determined by our exploration experiments. They can effectively present the characteristics of a task’s search space. Refer to Appendix D.1 for more details of our evaluation tasks.
Benchmark algorithms. In our experiment, we benchmark all tree-search algorithms introduced in section 3.2.2, including MCTS-$\alpha$, MCTS-Rollout, MCTS, BFS, and DFS. We mainly compare them with direct decoding methods: CoT greedy decoding (denoted as CoT), and CoT with self-consistency (denoted as CoT-SC). Considering the search space gap between direct decoding and tree decoding (especially the sentence-level action node discussed in the section 3.1), we also include another baseline - CoT-SC-Tree which conducts CoT-SC over the tree’s sentence nodes.
[…]
Model and training details. For the rollout policy used in tree-search, we use LLaMA2-7B (Touvron et al., 2023b) on three reasoning tasks, and GPT-2-small (125M) on the RLHF alignment task. All LLM will be first supervise-finetuned on the tasks’ training set, enabling their zero-shot CoT ability. Thus, CoT in our experiments refers to zero-shot greedy decoding. For value and ORM training, the data are generated by randomly sampling the SFT policy’s rollouts on the training set.
4.2 Results and Discussion
Question 1: How does TS-LLM perform in different generation tasks regarding Path@1? Does Path@1 provide a reasonable evaluation metric? (Sec. 3.2.2 and Sec. 3.4)
[…]
Method Performance(%) / # Tokens Reward / # Forward
GSM8k Game24 PrOntoQA RLHF(token-level)
CoT-greedy 41.4 $\pm$ 0.0 98 12.7 $\pm$ 0.0 76 48.8 $\pm$ 0.0 77 0.387 $\pm$ 0.0 57.8
MCTS-$\alpha$ 51.9 $\pm$ 0.6 561 63.3 $\pm$ 1.9 412 99.4 $\pm$ 0.2 190 2.221 $\pm$ 0.0 186
MCTS-Rollout 47.8 $\pm$ 0.8 3.4k 71.3 $\pm$ 2.5 670 96.9 $\pm$ 0.6 210 1.925 $\pm$ 0.0 809
MCTS 52.2 $\pm$ 1.7 486 64.0 $\pm$ 3.2 371 94.2 $\pm$ 1.6 125 -1.295 $\pm$ 0.0 61.8
BFS 52.5 $\pm$ 1.3 485 64.8 $\pm$ 2.9 369 94.4 $\pm$ 0.3 126 -1.295 $\pm$ 0.0 61.8
DFS 51.8 $\pm$ 0.6 486 66.3 $\pm$ 1.9 369 93.3 $\pm$ 0.8 126 -1.295 $\pm$ 0.0 61.8
[…]
Table 3: Performance on Game24 with different node size.
Table 4: Results of TS-LLM variants on Game24 by prompting LLaMA2-7B as the value function and the final step ORM.
Method Performance(%) / # tokens
expand by 6 expand by 20 expand by 50
MCTS-$\alpha$ 44.2 252 63.3 413 74.6 569
Rollout 44.2 346 71.3 672 79.8 838
BFS 45.9 202 64.8 373 74.9 527
CoT-SC-Tree@10 34.0 508 48.3 651 49.2 706
CoT-SC@10 - - - - 50.8 0.8k
Method Performance(%)
CoT-greedy 12.71 $\pm$ 0.0
BFS@1 9.2 $\pm$ 3.3
MCTS-$\alpha$@1 8.08 $\pm$ 0.3
Table 4: Results of TS-LLM variants on Game24 by prompting LLaMA2-7B as the value function and the final step ORM.
In Table 2, we first provide the comparison between TS-LLM variants with the CoT baseline regarding the Path@1 performance. Path@1 means each algorithm only generates one answer. TS-LLM variants generally outperform the baseline, underscoring the advantages of tree-search algorithms due to their capability to efficiently prune the search tree. However, notable exceptions are also seen in the RLHF task where MCTS, BFS, and DFS underperform. From our results, we want to highlight two things. Firstly, MCTS, BFS, and DFS generally perform better in shallow search problems (8 for GSM8K and 4 for Game24) while MCTS-$\alpha$ and MCTS-Rollout are dominant in deep search problems (15 for PrOntoQA and 64 for RLHF). Secondly, the the backward/backpropagation operation is quite important, since the first three methods—MCTS, BFS, and DFS—do not incorporate this operation and depend greedily on the value function. Note that the MCTS Path@1 also does not include the backward operation because it only happens after finding one complete path.
[…]
Question 2: How do different node constructions influence the performance? (Sec. 3.1)
As discussed in Sec. 3.1), the search space for sentence-level action nodes is limited. Thus, we investigate the possible influence introduced by different tree constructions on Game24 with different node sizes. Specifically, we set the number of maximal expanded node size as 6, 20, and 50. Table 4 lists the performance and the number of tokens generated comparing TS-LLM’s variants, CoT-SC and CoT-SC-Tree. The almost doubled performance boost from 44.2 to 79.8 indicates the importance of appropriate expansion node size, improving TS-LLM’s performance upper bound.
Question 3: Why do we need a learned value function and how to train that? (Sec. 3.2.1)
Table 5: Different value training for iterative update on GSM8k
Figure 2: Different Value training on GSM8k [Uncaptioned image]
Method Policy Value Performance(%)
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{0}}$ 51.9 $\pm$ 0.6
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{1}}^{\text{RL}}$ 52.0 $\pm$ 0.5
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{1}}$ 53.2 $\pm$ 0.3
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{0}}$ 54.1 $\pm$ 0.9
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{1}}^{\text{RL}}$ 55.2 $\pm$ 1.2
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{1}}$ 56.5 $\pm$ 0.6
Table 5: Different value training for iterative update on GSM8k
Figure 3: mean/max reward for RLHF alignment and the best results of 3 aggregations for the rest three tasks w.r.t. number of sequences on the 1st row and number of computations on the 2nd row.
In Table 4, we provide a motivating example by prompting LLaMA2-7b-base as the value function and reward model for TS-LLM on Game24. The performance drop of BFS and MCTS-$\alpha$ indicates the incapability of small language models as reliable evaluators, which motivates us to substitute it with a learned value function and ORM as a general solution for any task and LLM with any size.
[…]
Question 4: How TS-LLM is improved by aggregating over multiple results? (Sec. 3.2.3)
In Fig. 3, we demonstrate the mean/max reward for the RLHF task and the best of 3 aggregation results for the rest three tasks. We measure the performance of aggregation w.r.t path number and token consumption. From the figure, we mainly summarize two conclusions: Firstly, Most TS-LLM variants benefit from aggregation and can show large strengths compared with other baselines. CoT-SC only beats TS-LLM in GSM8k with the same token size, mainly because of its larger search space. TS-LLM variants are still dominant when compared with CoT-SC-Tree. Secondly, tree-search algorithms’ aggregation benefits less than CoT-SC in small-scale problems. In GSM8K and Game24, TS-LLM struggles to improve under large aggregation numbers. We believe this is because of: (1) The search space gap between CoT-SC and tree-search algorithms. Tree-search algorithms inherently explore fewer sentences, which is validated by comparing token consumption between CoT-SC-Tree@50 and CoT-SC@50. (2) Different tree searches are not independent. The latter search might be influenced by the previous one, which decreases generation diversity.
Question 5: Can TS-LLM further train LLM iteratively? (Sec. 3.3)
Table 6: Iterative update results. $\theta_{0}$ is the old parameter while $\theta_{1}$ is the new one after one iteration. We compare all combinations of policy and value on GSM8k (left) and RLHF alignment (right).
Method Policy Value Performance(%)
Greedy $\pi_{\theta_{0}}$ 41.4 $\pm$ 0.0
Greedy $\pi_{\theta_{1}}$ 47.9 $\pm$ 0.0
Greedy RFT k=50 - 47.0 $\pm$ 0.0
Greedy RFT k=100 - 47.5 $\pm$ 0.0
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{0}}$ 51.9 $\pm$ 0.6
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{1}}$ 53.2 $\pm$ 0.3
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{0}}$ 54.1 $\pm$ 0.9
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{1}}$ 56.5 $\pm$ 0.6
Method Policy Value Performance(%)
Greedy $\pi_{\theta_{0}}$ 0.39 $\pm$ 0.0
Greedy $\pi_{\theta_{1}}$ 1.87 $\pm$ 0.0
Greedy RFT N=5 - 1.16 $\pm$ 0.0
Greedy PPO - 2.53 $\pm$ 0.0
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{0}}$ 2.221 $\pm$ 0.0
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{1}}$ 2.482 $\pm$ 0.0
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{0}}$ 2.532 $\pm$ 0.0
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{1}}$ 2.670 $\pm$ 0.0
To answer this question, we conduct initial experiments for one iterative update. We utilize MCTS-$\alpha$ with old policy $\pi_{\theta_{0}}$, value $v_{\theta_{0}}$ and ORM $\hat{r}_{\theta_{0}}$, to sample answers on the training dataset as an augmentation to the origin one. It will be further used to finetune these models to ($\pi_{\theta_{1}},v_{\theta_{1}},\hat{r}_{\theta_{1}}$). We include two baselines, RFT (Yuan et al., 2023b), which utilizes rejection sampling to finetune the policy, with different sampling numbers $k$ or top $N$, and PPO (Schulman et al., 2017) for the RLHF task. Note that PPO conducts multiple iterative updates. It is not fully comparable to our method and we only add it for completeness. Refer to Appendix D.7 and D.8 for more experimental details.
In Table. 6, we list results of iterative update on the GSM8K and RLHF, covering greedy decoding and MCTS-$\alpha$ over all policy and value combinations. Our empirical results validate that TS-LLM can further train LLM policy, value and ORM, boosting performance with the new policy $\pi_{\theta_{1}}$, new value and ORM $\{v,\hat{r}\}_{\theta_{1}}$, or both $(\pi_{\theta_{1}},\{v,\hat{r}\}_{\theta_{1}})$ in CoT greedy decoding and MCTS-$\alpha$. $\pi_{\theta_{1}}$’s greedy performance is even slightly better than RFT which is specifically designed for GSM8k. We believe by further extending TS-LLM to multi-update, we can make it more competitive though currently $\pi_{\theta_{1}}$ still cannot beat PPO-based policy. For the training of value function and ORM (Question 3), we compare MCTS-$\alpha$ in Table 5. We train value and ORM in two paradigms, one ($\{v,\hat{r}\}_{\theta_{1}}$) is optimized from the initial weights and mixture of old and new tree-search data; another($\{v,\hat{r}\}_{\theta_{1}}^{\text{RL}}$) is optimized from $\{v,\hat{r}\}_{\theta_{0}}$ with only new tree-search data. This is called RL because training critic model in RL utilizes a similar process of continual training. The results show that $\{v,\hat{r}\}_{\theta_{1}}$ outperforms $\{v,\hat{r}\}_{\theta_{1}}^{\text{RL}}$ on both old and new policy when conducting tree search, contrary to the normal situation in traditional iterative RL training.
5 Conclusion
In this work, we propose TS-LLM, an LLM inference and training framework guided by Alphazero-like tree search that is generally versatile for different tasks and scaled to token-level expanded tree spaces. Empirical results validate that TS-LLM can enhance LLMs decoding and serve as a new training paradigm. We leave further discussion of our limitation and future work in Appendix B.
Appendix A More Related Work and Comparisons
Here we discuss the differences between TS-LLM and relevant work mentioned in Sec 2 in detail.
Recent efforts have been made to investigate non-linear reasoning structures such as trees(Jung et al., 2022; Zhu et al., 2022; Xu, 2023; Xie et al., 2023; Yao et al., 2023; Hao et al., 2023). Approaches for searching on trees with LLM’s self-evaluation have been applied to find better reasoning paths, e.g. beam search in Xie et al. (2023), depth-/breadth-first search in Yao et al. (2023) and Monte-Carlo Tree Search in (Hao et al., 2023). Compared with these methods, TS-LLM is a tree search guided LLM decoding framework with a learned value function, which is more generally applicable to reasoning tasks and other scenarios like RLHF alignment. TS-LLM includes comparisons between different tree search approaches, analysis of computation cost, and shows the possibility of improving both the language model and value function iteratively. The most relevant work, CoRe (Zhu et al., 2022), proposes to finetune both the reasoning step generator and learned verifier for solving math word problems using MCTS for reasoning decoding which is the most relevant work to ours. Compared with CoRe, in this work TS-LLM distinguishes itself by:
1. TS-LLM is generally applicable to a wide range of reasoning tasks and text generation tasks under general reward settings, from sentence-level trees to token-level trees. But CoRe is proposed for Math Word Problems and only assumes a binary verifier (reward model).
2. In this work, we conduct comprehensive comparisons between popular tree search approaches on reasoning, planning, and RLHF alignment tasks. We fairly compare linear decoding approaches like CoT and CoT-SC with tree search approaches w.r.t. computation efficiency.
3. TS-LLM demonstrates potentials to improve LLMs’ performance of direct decoding as well as tree search guided decoding, while in CoRe the latter cannot be improved when combining the updated generator (language model policy) with the updated verifier (value function) together.
Other related topic:
Search guided decoding in LLMs Heuristic search and planning like beam search or MCTS have also been used in NLP tasks including machine translation (Leblond et al., 2021) and code generation (Zhang et al., 2023; Matulewicz, 2022). During our preparation for the appendix, we find a concurrent work (Liu et al., 2023) which is proposed to guide LLM’s decoding by reusing the critic model during PPO optimization to improve language models in alignment tasks. Compared with this work, TS-LLM focuses on optimizing the policy and value model through tree search guided inference and demonstrates the potential of continuously improving the policy and value models. And TS-LLM is generally applicable to both alignment tasks and reasoning tasks by conducting search on token-level actions and sentence-level actions.
Appendix B Limitation and future work
Currently, our method TS-LLM still cannot scale to really large-scale scenarios due to the extra computation burdens introduced by node expansion and value evaluation. Additional engineering work such as key value caching is required to accelerate the tree-search. In addition, we do not cover all feasible action-space designs for tree search and it is flexible to propose advanced algorithms to automatically construct a tree mixed with both sentence-level expansion and token-level expansion, etc. We leave such exploration for future work. For MCTS aggregation, the current method still struggles to improve under large aggregation numbers. some new algorithms that can encourage multi-search diversity might be needed. Currently, we are still actively working on scaling up our method both during inference and training (especially multi-iteration training).
Appendix C Background of Monte Carlo Tree-search Algorihtms
Once we build the tree, we can use various search algorithms to find a high-reward trace. However, it’s not easy to balance between exploration and exploitation during the search process, especially when the tree is sufficiently deep. Therefore we adopt Monte Carlo Tree Search(MCTS) variants as choices for strategic and principled search. Instead of the four operations in traditional MCTS (Kocsis & Szepesvári, 2006; Coulom, 2006), we refer to the search process in AlphaZero (Silver et al., 2017a) and introduce 3 basic operations of a standard search simulation in it as follows, when searching actions from current state node $s_{0}$:
Select It begins at the root node of the search tree, of the current state, $s_{0}$, and finishes when reaching a leaf node $s_{L}$ at timestep $L$. At each of these $L$ timesteps(internal nodes), an action(child node) is selected according to $a_{t}=\displaystyle\arg\max\limits_{a}\left(Q(s_{t},a)+U(s_{t},a)\right)$ where $U(s_{t},a)$ is calculated by a variant of PUCT algorithm (Rosin, 2011):
$U(s,a)=c_{\text{puct}}\cdot\pi_{\theta}(s,a)\frac{\sqrt{\sum_{b}N(s,b)}}{1+N(s,a)}$ (2)
where $N(s,a)$ is the visit count of selecting action $a$ at node $s$, and $c_{\text{puct}}=\log((\sum_{b}N(s,b)+c_{\text{base}}+1)/c_{\text{base}})+c_{\text{init}}$ is controlled by visit count and two constants. This search control strategy initially prefers actions with high prior probability and low visit count, but asymptotically prefers actions with high action-value.
Expand and evaluate After encountering a leaf node $s_{L}$ by select, if $s_{L}$ is not a terminal node, it will be expanded by the language model policy. The state of the leaf node is evaluated by the value network, noted as $v(s_{L})$. If $s_{L}$ is a terminal node, if there is an oracle reward function $R$, then $v(s_{L})=R(s_{L})$, otherwise, in this paper, we use an ORM $\hat{r}$ as an approximation of it.
Backup After expand and evaluate on a leaf node, backward the statistics through the path $s_{L},s_{L-1},\ldots,s_{0}$, for each node, increase the visit count by $N(s_{t},a_{t})=N(s_{t},a_{t})+1$, and the total action-value are updated as $W(s_{t},a_{t})=W(s_{t},a_{t})+v(s_{L})$, the mean action-value are updated as $Q(s_{t},a_{t})=W(s_{t},a_{t})/N(s_{t},a_{t})$.
In this paper, we introduce 3 variants of MCTS based on the above basic operations.
Appendix D Experiment Details
D.1 Task setups
GSM8k GSM8k (Cobbe et al., 2021) is a commonly used numerical reasoning dataset, Given a context description and a question, it takes steps of mathematical reasoning and computation to arrive at a final answer. There are about 7.5k problems in the training dataset and 1.3k problems in the test dataset.
[…]
PrOntoQA PrOntoQA (Saparov & He, 2022) is a typical logical reasoning task in which a language model is required to verify whether a hypothesis is true or false given a set of facts and logical rules. There are 4k problems in the training dataset and 500 problems in the test dataset.
RLHF We choose a synthetic RLHF dataset Dahoas  [^1] https://huggingface.co/datasets/Dahoas/synthetic-instruct-gptj-pairwise serving as the query data. We split the dataset to 30000/3000 as training and test set respectively. For the reward model, we choose reward-model-deberta-v3-large-v2222https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2 from OpenAssistant, which is trained from several RLHF datasets.
D.2 SFT and value training details
SFT in GSM8k, Game24 and PrOntoQA: For GSM8k, Game24 and PrOntoQA, we first train LLaMA2-7b on the training dataset The training is conducted on 8 NVIDIA A800 GPUs, using a cosine scheduler decaying from lr=2e-5 to 0.0 with a warmup ratio of 0.03, batch size 128 for 3 epochs. For GSM8k and Game24 we use the checkpoint at the last epoch as the direct policy in experiments, while for PrOntoQA we use the checkpoint at the 1st epoch since the others overfit.
Value training in GSM8k, Game24 and PrOntoQA: Then we train the value function on the data rollout by the SFT policy. In GSM8k and Game24, For each model checkpoints of 3 epochs during SFT, we first collect 100 outputs per problem in the training dataset, then duplicate the overlapped answers, labeled each answer with our training set outcome reward ocracle. For data sampled by ech model checkpoint, we subsample 17 answers per problem, which is in total at most 51 answers per problem after deduplication. In PrOntoQA, we only sample 50 answers per problem with the first epoch model checkpoint and then do deduplication.
The value functions are trained in the same setting as supervised finetuning. We set the reward to be 1 when the output answer is correct and -1 otherwise. Then we use MC with $\gamma=1$ to compute the returns. We do model selection on a validation dataset sampled from the direct policy model. For GSM8k, we train the value function and ORM for one epoch, while for Game24 and PrOntoQA we train the value function and ORM for 3 epochs.
SFT in RLHF alignment: We utilize GPT2-open-instruct [^3] https://huggingface.co/vicgalle/gpt2-open-instruct-v1, a GPT2-Small model supervised-finetuned over several instruction-tuning dataset.
Value training in RLHF alignment: Based on the SFT model, we collect 50 rollouts by the SFT policy for each question in the training set and label their final reward with the reward model. Then we train the value function and ORM for 2 epochs.
Note that here we start training the value function and ORM from the data sampled by the SFT policy model through direct decoding just as an initialization of the value function and ORM. After that TS-LLM can optimize the policy model, value function, and ORM simultaneously by adding new data sampled from tree search into the training buffer.
D.3 Details of value dataset ablation
Here we introduce the details of building mixed, pure and pure,less datasets on GSM8k for value training in Fig 2. For each model checkpoints of 3 epochs during SFT, we first collect 100 outputs per problem in GSM8k training dataset, and then duplicate the overlapped answers, labeled each answer with our training set outcome reward ocracle. we sample multiple output sequences with temperature=0.7, top p=1.0 and top k=100.
mixed dataset: For each deduplicated dataset sampled by models of 3 epochs, we subsample 17 answers per problem.
pure dataset: we subsample 50 answers per problem from deduplicated dataset sampled by the last epoch policy model.
pure,less dataset: we subsample 17 answers per problem from deduplicated dataset sampled by the last epoch policy model.
For the results in Table 5, the details of training $\{v,\hat{r}\}_{\theta_{1}}$ can be find in Sec D.8. We use MC with $\gamma=1$ to compute the returns. Here we describe the details of training $\{v,\hat{r}\}_{\theta_{1}}^{\text{RL}}$, we use the collected 78.7k samples in Sec D.7 to optimize $\{v,\hat{r}\}_{\theta_{0}}$. The training uses a cosine scheduler decaying from lr=2e-5 to 0.0 with a warmup ratio of 0.03, batch size 128 for 3 epochs.
D.4 Details of applying each tree search approach
We present the implementation details and hyperparameters of all tree search approaches here.
Firstly, we refer to Table 1 for basic settings of each task. We set temperature=1.0, top_p=1.0, top_k=100 when using LLM to generate tree actions. To compute logprobs when expand actions in RLHF alignment task trees, we also use a temperature of 1.0.
For MCTS variants including MCTS-$\alpha$, MCTS-Rollout and MCTS, we need to define the hyperparamter in PUCT algorithm:
$c_{\text{puct}}=\log((\sum_{b}N(s,b)+c_{\text{base}}+1)/c_{\text{base}})+c_{\text{init}}$
In this paper, we fixed $c_{\text{base}}$ with 19652 and set $c_{\text{init}}=3$ for GSM8k, Game24 and RLHF alignment tasks, set $c_{\text{init}}=0.3$ for PrOntoQA tasks. Specifically, in MCTS-$\alpha$, we set the number of simulations before making an action to 5 for GSM8k, Game24 and PrOntoQA, 10 for RLHF alignment. We deterministically sampled actions with the largest visit count. And in MCTS-Rollout, we set an computation upperbound as number of generated tokens or number of model forwards, which is 51200 in GSM8k and Game24, 1800 in PrOntoQA and 5000 in RLHF alignment.
For depth-first search (DFS), the children of a non-leaf node is traversed in a non-decreasing order by value. For efficient exploration, we tried 2 heuristics to prune the subtrees, (1) drop the children nodes with low value by prune_ratio. (2) drop the children nodes lower than a prune_value. We set prune_ratio to be 0.7 for GSM8k, Game24 and PrOntoQA, and 0.9 for RLHF alignment task.
All Path@1 results for each tree search approach is conducted with 3 seeds and show the mean and standard deviation. Note that for Path@1 results of tree search approaches, the randomness comes from the node expansion process where we use an LLM to sample candidate actions. While for CoT-SC results, the randomness comes from sampling during direct decoding.
D.5 Details of aggregation results
Here we present the details of how we sample multiple answers with tree search approaches and aggregate them into a final answer.
When sampling multiple intra-tree answers with MCTS-$\alpha$, we use a stochastic sampling setting. To ensure MCTS-$\alpha$ to explore sufficently, when selecting action of the current node, before doing several times of simulation, we add Dirichlet noise into the language model’s prior probability of the current root node $s_{0}$, i.e. $\pi^{\prime}_{\theta}(s_{0},a)=(1-\epsilon)\pi_{\theta}(s_{0},a)+\epsilon\eta$, where $\eta\sim\text{Dir}(0.3)$, and we set $\epsilon=0.25$ for both tasks. Actions are sample based on visit count, i.e. $a\sim\frac{N(s_{t},a)^{1/\tau}}{\sum_{b}N(s_{t},b)^{1/\tau}}$, where we set $\tau=1$. After returning with a complete path, we clear the node statistics ($Q(s_{t},a_{t})$ and $N(s_{t},a_{t})$) on the tree to eliminate the influence of previous searches, while the tree structure is maintained. This setting is denoted as clear-tree when presented in the table. In summary, when we only measure path@1 performance, we adopt MCTS-$\alpha$ (no sampling). But when we measure the aggregation performance, we use MCTS-$\alpha$-intra tree or MCTS-$\alpha$-inter tree. In MCTS-$\alpha$-intra tree we will activate the clear-tree and stochastic sampling setting.
When sampling multiple answers with other tree-search methods, we only utilize the intra-tree aggregation variant, without stochastic sampling and clear-tree setting. This is because only MCTS-$\alpha$ and MCTS-rollout can conduct the above sampling and clear-tree operation. And we temporarily only apply such setting on MCTS-$\alpha$.
In ORM-vote, since we train the ORM with the reward signal -1 and 1, given a list of $N$ answers to be aggregated, we first normalize its values $\{\hat{r}(y^{k})\}_{k=0}^{N-1}$ with min-max normalization make them in $[0,1]$.
D.6 Results per task per aggregation
We show detailed results of GSM8k on Table 7, results of Game24 on Table 8, Table 9 for PrOnToQA and Table 10 for RLHF alignment.
Table 7: Detailed Results in GSM8k
Method N Majority-vote ORM-vote ORM-max #Token
CoT - 41.4 41.4 41.4 0.1k
CoT-SC 1 38.21 38.21 38.21 0.1k
CoT-SC 10 51.93 57.47 53.83 1k
CoT-SC 20 54.44 59.44 54.74 2k
CoT-SC 50 56.79 61.03 54.44 5k
CoT-SC 100 58.15 62.70 53.68 10k
CoT-SC-Tree 1 37.91 37.91 37.91 0.1k
CoT-SC-Tree 10 50.19 53.15 50.95 0.8k
CoT-SC-Tree 20 52.69 55.12 52.84 1.3k
CoT-SC-Tree 50 54.51 57.16 53.45 2.7k
BFS 1 52.5 52.5 52.5 0.5k
BFS 10 58.98 56.25 54.97 3.1k
BFS 20 58.91 56.79 52.39 5.3k
BFS 50 59.29 59.36 53.22 10.1k
DFS 1 51.8 51.8 51.8 0.5k
DFS 10 53.75 54.66 53.37 1.2k
DFS 20 54.89 55.50 53.53 1.6k
MCTS-$\alpha$ (no sampling) 1 51.9 51.9 51.9 0.5k
MCTS-$\alpha$-intra tree 1 46.78 46.78 46.78 0.7k
MCTS-$\alpha$-intra tree 10 57.85 56.86 54.36 3.4k
MCTS-$\alpha$-intra tree 20 58.83 58.23 55.19 5.3k
MCTS-$\alpha$-inter trees 1 51.9 51.9 51.9 0.5k
MCTS-$\alpha$-inter trees 10 57.92 58.53 55.34 5.5k
MCTS-$\alpha$-inter trees 20 58.83 59.06 54.97 11.1k
MCTS-$\alpha$-inter trees 50 58.76 61.26 53.98 27.8k
MCTS 1 52.2 52.2 52.2 0.5k
MCTS 10 57.92 55.72 53.75 2.4k
MCTS 20 58.61 56.79 54.74 4.0k
MCTS 50 59.36 58.23 53.75 7.5k
MCTS-Rollout 1 47.8 47.8 47.8 3.4k
MCTS-Rollout 10 51.10 50.49 49.81, 5.4k
MCTS-Rollout 20 51.86 51.25 50.19 6.1k
MCTS-Rollout 50 52.69 52.24 50.49 7.2k
Table 8: Detailed Results in Game24
Method N Majority-vote ORM-vote ORM-max #Token
CoT - 12.7 12.7 12.7 0.1k
CoT-SC 1 9.94 9.94 9.94 0.1k
CoT-SC 10 13.54 50.83 50.83 0.8k
CoT-SC 20 14.36 65.75 65.47 1.6k
CoT-SC 50 16.30 78.45 78.45 4.0k
CoT-SC 100 18.23 84.25 84.53 7.9k
CoT-SC-Tree 1 9.67 9.67 9.67 0.1k
CoT-SC-Tree 10 11.33 48.34 48.34 0.7k
CoT-SC-Tree 20 13.26 61.60 62.15 1.1k
CoT-SC-Tree 50 16.57 69.61 69.89 2.0k
BFS 1 64.8 64.8 64.8 0.4k
BFS 10 47.79 70.72 70.99 1.6k
BFS 20 27.62 69.34 69.34 2.3k
BFS 50 7.18 70.17 70.72 3.7k
DFS 1 66.3 66.3 66.3 0.4k
DFS 10 50.83 62.98 64.09 0.4k
DFS 20 50.83 62.98 64.09 0.6k
MCTS-$\alpha$ (no sampling) 1 63.3 63.3 63.3 0.4k
MCTS-$\alpha$-intra tree 1 64.36 64.36 64.36 0.4k
MCTS-$\alpha$-intra tree 10 66.85 67.68 63.90 0.9k
MCTS-$\alpha$-intra tree 20 67.13 69.34 68.78 1.1k
MCTS-$\alpha$-intra tree 50 67.96 69.89 69.34 1.4k
MCTS-$\alpha$-inter trees 1 63.3 63.3 63.3 0.4k
MCTS-$\alpha$-inter trees 10 72.65 82.87 82.32 4.1k
MCTS-$\alpha$-inter trees 20 72.93 84.25 83.15 8.3k
MCTS 1 64.0 64.0 64.0 0.4k
MCTS 10 70.44 70.72 70.17 0.8k
MCTS 20 72.10 72.10 71.27 1.1k
MCTS 50 72.38 72.38 71.55 1.6k
MCTS-Rollout 1 71.3 71.3 71.3 0.7k
MCTS-Rollout 10 73.48 73.20 72.65 0.9k
MCTS-Rollout 20 73.48 73.48 72.38 1.0k
MCTS-Rollout 50 73.48 73.48 72.38 1.1k
Table 9: Detailed Results in PrOntoQA
Method N Majority-vote ORM-vote ORM-max #Token
CoT - 48.8 48.8 48.8 77
CoT-SC 1 54.40 54.40 54.40 91.25
CoT-SC 3 63.60 82.40 82.40 273.75
CoT-SC 10 58.40 97.80 97.80 912.55
CoT-SC 20 57.00 99.80 99.80 1.8k
CoT-SC-Tree 1 50.20 50.20 50.20 82.02
CoT-SC-Tree 10 62.40 98.40 98.40 413.58
CoT-SC-Tree 20 61.00 99.40 99.40 632.91
BFS 1 94.40 94.40 94.40 125.52
BFS 10 99.00 100.00 100.00 837.78
BFS 20 98.60 99.80 99.80 1.5k
DFS 1 93.30 93.30 93.30 124.46
DFS 10 95.60 96.20 96.00 178.79
DFS 20 95.60 96.20 96.20 180.92
MCTS-$\alpha$ (no sampling) 1 99.40 99.40 99.40 183.66
MCTS-$\alpha$-intra tree 1 97.20 97.20 97.20 208.68
MCTS-$\alpha$-intra tree 10 99.80 99.80 99.80 364.96
MCTS-$\alpha$-intra tree 20 99.80 99.80 99.80 441.31
MCTS-$\alpha$-inter trees 1 99.40 99.40 99.40 183.66
MCTS-$\alpha$-inter trees 10 100.00 100.00 100.00 1.9k
MCTS-$\alpha$-inter trees 20 100.00 100.00 100.00 3.8k
MCTS 1 94.20 94.20 94.20 126.65
MCTS 10 99.60 99.60 99.60 182.88
MCTS 20 100.00 100.00 100.00 240.16
MCTS-Rollout 1 96.90 96.90 96.90 210.41
MCTS-Rollout 10 99.20 99.20 99.20 220.16
MCTS-Rollout 20 99.20 99.20 99.20 224.16
Table 10: Detailed Results in RLHF alignment
Method N Mean Best #Forward
CoT 1 0.387 0.387 57.8
CoT-SC 1 -0.164 -0.164 58
CoT-SC 10 -0.182 1.592 0.6k
CoT-SC 20 -0.175 1.972 1.2k
CoT-SC 50 -0.176 2.411 2.9k
BFS 1 -1.295 -1.295 61.8
BFS 10 -1.523 -1.065 0.6k
BFS 20 -1.520 -0.948 1.2k
BFS 50 -1.474 -0.813 3.1k
DFS 1 -1.295 -1.295 61.8
DFS 10 -1.67 -1.17 64.8
DFS 20 -1.69 -1.08 66.9
DFS 50 -1.71 -0.99 71.9
MCTS-$\alpha$ (no sampling) 1 2.221 2.221 186
MCTS-$\alpha$-intra tree 1 1.538 1.538 198.50
MCTS-$\alpha$-intra tree 10 1.527 3.052 1.6k
MCTS-$\alpha$-intra tree 20 1.533 3.311 3.1k
MCTS 1 -1.295 -1.295 61.8
MCTS 10 -1.146 0.160 0.6k
MCTS 20 -1.08 0.528 1.2k
MCTS 50 -0.961 0.981 2.8k
MCTS-Rollout 1 1.925 1.925 0.8k
MCTS-Rollout 10 2.278 2.540 1.1k
MCTS-Rollout 20 2.376 2.643 1.2k
MCTS-Rollout 50 2.491 2.746 1.3k
D.7 sampling details of iterative update
We verify the idea of iteratively enhancing language model policy and value function model on the GSM8k and RLHF datasets.
Sampling in GSM8k: When sampling from the 7.5k problems in the GSM8k training dataset, we sample 12 sequences per problem in one sentence-level expanded tree, after deduplication, this results in 78.7k distinct answers, and 73.2% are correct answers. The sample parameters are listed in Table 11.
Sampling in RLHF alignment: We collect 10 answers for each training set problem sampled by MCTS-$\alpha$. We list the specific hyperparameters in Table 12.
To collect data for the rejection sampling baseline, we first sample 10 sequences per problem and then use the top 5 sequences for supervised fine-tuning.
Table 11: Hyperparameters of sampling in GSM8k for LLM decoding(left), tree construction setting(middle), and MCTS-$\alpha$ setting(right).
Hyperparameter value
temperature 1.0
top_p 1.0
top_k 100
Hyperparameter value
Tree Max width 6
Tree Max depth 8
Node Sentence
Hyperparameter value
num simulation 5
clear tree True
stochastic sampling True
$c_{\text{base}}$ 19652
$c_{\text{init}}$ 3
$\tau$ 1.0
Table 12: Hyperparameters of sampling in RLHF alignment for LLM decoding(left), tree construction setting(middle), and MCTS-$\alpha$ setting(right).
Hyperparameter value
temperature 1.0
top_p 1.0
top_k 50
Hyperparameter value
Tree Max width 50
Tree Max depth 64
Node Token
Hyperparameter value
num simulation 5
clear tree True
stochastic sampling True
$c_{\text{base}}$ 19652
$c_{\text{init}}$ 3
$\tau$ 1.0
D.8 Training details of iterative update
Policy training in GSM8k: We construct the dataset for supervised finetuning by combining data in the training dataset with 57.6k correct answers sampled in Sec D.7 which results in 64.1k distinct correct answers. And we train the new policy model $\pi_{\theta_{1}}$ from the starting base model LLaMA2-7b for 3 epochs, following Yuan et al. (2023a). The training setting is the same as described in Sec D.2.
Value training in GSM8k: We construct the dataset for value and ORM training by combining the data used to train $\{v,\hat{r}\}_{\theta_{0}}$ with 78.7k answers sampled by MCTS-$\alpha$ in Sec D.7. To fairly compare $\{v,\hat{r}\}_{\theta_{1}}$ with $\{v,\hat{r}\}_{\theta_{0}}$, we drop samples in the former dataset to keep at most $51-12=39$ answers per problem resulting in 359k distinct answers. And we train the new value function $\{v,\hat{r}\}_{\theta_{1}}$ from the value model with its initial weight(before being updated on any data) for 3 epochs. The training setting is the same as described in Sec D.2.
Policy training in RLHF alignment: For the MCTS-$\alpha$’s training, we subsample the top 5 answers from the full 10 candidates (mentioned in Appendix D.7) to serve as the SFT dataset. For the RFT n=5 baseline, we subsample the top 5 answers from 50 direct decodings as the SFT dataset. For the training of the PPO algorithm, we adopt the implementation from trlx [^4] https://github.com/CarperAI/trlx. We sample 20 answers for each question in total, which maintains the same level of token consumption during the PPO rollouts as that of MCTS-$\alpha$.
Value training in RLHF alignment: We construct the value and ORM dataset by mixing data from SFT-policy direct decoding and from MCTS-$\alpha$. To make the comparison fair, the new value function’s training utilizes the same amount of data as the old one by subsampling 40 answers (from 50 shown in Appendix D.2) from direct decoding data and all 10 answers (shown in Appendix D.7) generated by MCTS-$\alpha$. We train our value function with learning rate 2e-5 and cosine scheduler from the initial model (instead of continuing training from the old value function) for 2 epochs.
Appendix E Qualitative Results
Here we present some qualitative results of generation for different methods (Path@1 performance) in Table 13 (for Game24) and 14 (for the RLHF alignment task).
Prompt: 4 8 9 13
Methods
Response
Correctness
CoT
8 + 13 = 21 (left: 4 9 21)
4 - 9 = -5 (left: -5 21)
21 - -5 = 26 (left: 26)
The answer is (8 + 13) - (4 - 9) = 26
No
BFS
9 + 3 = 12 (left: 4 8 12)
12 / 4 = 3 (left: 3 8)
3 * 8 = 24 (left: 24)
The answer is (9 + 3) / 4 * 8 = 24
No
MCTS-$\alpha$
13 - 9 = 4 (left: 4 4 8)
4 * 4 = 16 (left: 8 16)
16 + 8 = 24 (left: 24)
The answer is 4 * (13 - 9) + 8 = 24
Yes
MCTS-Rollout
13 - 9 = 4 (left: 4 4 8)
4 * 4 = 16 (left: 8 16)
16 + 8 = 24 (left: 24)
The answer is (13 - 9) * 4 + 8 = 24
Yes
Table 13: Qualitative examples of Game24 outputs
Prompt: What careers am I good at.
Methods
Response
Reward
CoT
You can be a great career at any career you choose. Depending on your interests and skills, you may be able to find a variety of jobs that align with your interests and skills.
-2.625
BFS
Some careers great careers to do professionally are entrepreneur in healthcare policy policy advocacy consulting public diplomacy leadership policy innovation in global issues managing information theory. Some great careers to achieve excellence in healthcare management include doctor psychologist
-2.03
MCTS-$\alpha$
You may be good at a variety of careers, depending on your interests and skills. Some of the most popular careers include software engineer, data scientist. You may also be good at a variety of fields such as finance, marketing, and business administration.
-1.67
Table 14: Qualitative examples of RLHF outputs